OpenStudy (anonymous):

The distance travelled by Dalton and Austin is in ratio 3:4. The distance travelled by Dalton is 18 miles. Find the distance travelled by Austin

OpenStudy (texaschic101):

3/4 = 18/x cross multiply 3x = (4)(18) 3x = 72 x = 72/3 x = 24 Austin traveled 24 miles.

OpenStudy (anonymous):

Ahan thanks :)

OpenStudy (texaschic101):

no problem :)